How many moles of sodium hypobromite (NaBrO) should be added to 1.00 L of 5.5×10^?2 M hypobromous acid (HBrO) to form a buffer solution of pH= 9.17? Assume that no volume change occurs when the NaBrO is added. I posted this question previously but it was answered incorrectly

How many moles of sodium hypobromite (NaBrO) should be added to 1.00 L of 5.5*10-2 M hypobromous acid (HBrO) to form a buffer solution of pH=9.17? Assume that no volume change occurs when the NaBrO is added Express your answer using two significant figures. 17 Azo ? mol
Answer #1

Formula of buffer for weak acid and strong base is

pH= pKa + log[salt]/[acid]

pKa for hypobromous acid= 8.65 since its Ka value=2.8×10-9

Now, putting the values pH=9.17, pka=8.65 , [HBrO]=5.5×10-2M

9.17=8.65+log[NaBrO]/[HBrO]

0.52=log[NaBrO]/[HBrO]

3.311=[NaBrO]/[HBrO]

[NaBrO]=0.1821 moles